Interactive Solution 6.39 presents a model for solving this problem. A slingshot fires a pebble from the top of a building at a speed of 14.0 m/s. The building is 32.0 m tall. Ignoring air resistance, find the speed with which the pebble strikes the ground when the pebble is fired (a) horizontally, (b) vertically straight up, and (c) vertically straight down.
Answer #1

By energy conservation

(1/2) mu2 + mgh = (1/2) m v2 + 0

v2 = u2 + 2gh = 142 + (2×9.8×32) = 823.2

v = 28.7 m/s

Answer will be same in all cases independent upon direction of projection.